### v1

[…]
H1.
Most of its components are 3D-printed using polylactic acid (PLA), leading to low rigidity and susceptibility to structural deformation.
H2.
Unlike typical robots with constrained ranges, its joints can rotate beyond 360∘, potentially reaching positions the robot cannot achieve.
H3.
Although portable, inevitable body movements during operations can cause its base to shift, leading to inaccurate action recording.
H4.
Its gripper lacks smooth control, leading to potential jamming under clamping forces.
H5.
Its shaft connects directly to the encoder, with wires routed along its side. Joint movement can cause constant friction and stretching, risking wire breakage or short circuits over long-time use.
During demonstration collection using teleoperation, most of the above issues (except H4) might not significantly impact data acquisition, as the human operator can adjust their actions based on the movements of the robotic arms. However, during the in-the-wild demonstration collection, these drawbacks can substantially affect motion capture accuracy and may result in invalid or unusable demonstration data. The hardware design of AirExo-2 is primarily driven by the limitations mentioned above. To ensure seamless integration with the learning process from in-the-wild demonstrations, we design the exoskeleton to match the dimensions of the robotic arm in a 1:1 ratio. This design choice helps avoid unnecessary obstacles in direct learning from the collected demonstrations. The key hardware designs are outlined as follows. Enhanced Overall Structural Rigidity (H1). In AirExo, the links connecting two consecutive joints are the most prone to deformation. In AirExo-2, we replace the 3D-printed parts with 20x20 European standard aluminum profiles, providing significantly higher strength at a very low cost. For the joints, the outer shell is 3D-printed using PLA-CF, a carbon fiber-reinforced PLA material with higher hardness. Inside the joint, larger bearings are used to further enhance structural rigidity. Together with the improved links, the hardware upgrade significantly increases the overall structural rigidity of AirExo-2, making the exoskeleton more durable, and thereby improving data collection accuracy. Hollow Rotating Disc and Side-Mounted Encoder (H5). As shown in Fig. H1., this design features an encoder mounted on the side of the joint, which uses a gear mechanism to translate the rotational angle of the rotating disc into encoder readings. The hollow disc design allows wires to pass through, preventing them from stretching during rotation and thereby extending their lifespan. Additionally, the side-mounted encoder simplifies maintenance, enabling easy debugging and replacement without the need to disassemble the joint. Joint with Angle Limit and Adjustable Friction (H2). This structure consists of a rotating disc with a grooved track and a friction pad that can be embedded into the track. Together, they allow for adjustment of the limiting angle, ensuring that the motion range of the AirExo-2 joint exactly aligns with the corresponding joint range of the robot. As illustrated in Fig. H1., the rotating disc, pre-joint, and post-joint are connected through bearings, allowing the rotational motion of the joint to be directly transmitted to the track. For demonstration collection, excessive or insufficient friction in the rotation of the joints is undesirable. Hence, the friction force of the joint in AirExo-2 can be adjusted by turning the screw on the outer shell, which compresses the friction pad. This design ensures optimal friction for comfortable and accurate data collection.
[width=]images/hardware.pdf
Smooth Gripper Control (H4). Following [umi, forcemimic], the gripper of AirExo-2 incorporates a linear guide, with the fingers mounted on a sliding block. This design allows for smoother opening and closing of the gripper, ensuring it operates seamlessly even under significant clamping forces without any stalling. Mobile Data Collection Platform (H3). Portability is crucial for in-the-wild data collection. However, to address the issue of base movement caused by the body motion of the operator, we mount AirExo-2 on a mobile aluminum profile stand, as shown in Fig. H1.. This setup ensures stability of the base during demonstration collection while maintaining the flexibility needed for mobility, enabling large-scale demonstration collection in real-world environments. An Intel RealSense D415 camera is set up on the top of the mobile platform to capture global observations. We also designed two optional camera mounts (though not used in this paper) for the future integration of in-hand cameras on the top of both grippers.
[…]
The AirExo-2 system requires two types of calibration simultaneously: (1) aligning the zero positions of each joint with the corresponding robot joint, and (2) determining the transformation between the global camera and the AirExo-2 base. To address these challenges, we propose a two-stage calibration process. Initial Calibration. For initial calibration, the former calibration can be achieved by manually adjusting the joints to approximate the zero position using specialized 3D-printed tools and reading the encoder values, obtaining . The latter calibration can be done by attaching an ArUco calibration marker board with a known position on the base and performing optical calibration using the OpenCV library [opencv], obtaining . Thus, the transformation can calculated as
However, this approach introduces errors due to human observation, calibration board misalignment, and optical inaccuracies. In a chained system like AirExo-2, these errors may propagate and amplify across joints, leading to significant cumulative inaccuracies in the end-effector pose. Therefore, fine-grained calibration is essential to ensure precise and consistent alignment between AirExo-2 and the camera frames during demonstration collection. Calibration via Differentiable Rendering. In the second stage, inspired by prior works [easyhec, easyhec++, sam-rl], we use differentiable rendering [differentiable_rendering] to refine the initial calibration. Training samples are obtained from a single human play trajectory with AirExo-2. Using the joint states and calibration parameters, we render the system mask and depth via a differentiable rendering engine [redner]. Calibration parameters are optimized by minimizing discrepancies between the rendered and annotated system masks, as well as between the rendered and observed depths. Pseudo-ground-truth masks are manually annotated with SAM-2 [sam2]. This iterative refinement compensates for errors that accumulate across joints, ultimately improving the overall accuracy. Specifically, we define , the calibration parameters to be optimized through differentiable rendering, as
where parameters, except for the base-to-camera rotation, are represented as deltas relative to the initial calibration results. The base-to-camera rotation is expressed in a 6D format [rot6d]. Thus, the final calibration results can be calculated as
{gather}
[…]
and the initial parameter values are set as
For the optimization process, we first record a single in-the-wild play trajectory, in which the human operator uses the AirExo-2 to adopt various poses. During trajectory recording, ensure that all parts of the AirExo-2 remain above the human hands and arms from the camera’s perspective. After data collection, we sample approximately 40 image-joint pairs from the trajectory, denoted as , where represents the total number of training samples for calibration, and and are the RGB and depth images of the -th sample, respectively. Subsequently, we utilize SAM-2 [sam2] to annotate the AirExo-2 mask and the depth mask for each sample , as shown in Fig. H1.. The first mask provides supervision for the rendered AirExo-2 mask, while the second mask is used to select the valid depth information that serves as the supervision signal.
[width=]images/calib_detail.pdf
The differentiable rendering engine Redner [redner] is employed to render the AirExo-2 mask and AirExo-2 depth using the calibration results and joint information:
where the rendering engine computes the gradients of the calibration parameters during the rendering process, and represent the initial calibration results. The rendered AirExo-2 mask is supervised by the human-annotated pseudo-AirExo-2 mask , and the rendered AirExo-2 depth is supervised by the camera depth within the region of the human-annotated depth mask . The depth mask ensures that only accurate depth information contributes to the loss. Thus, the objective can be written as:
where represents the weighting coefficient, and denotes the mask-apply operation. In practice, we set , use samples for optimization, and employ the Adam optimizer [adam] with a learning rate of for 1000 iterations to fine-tune the calibration results.
[…]
Apart from the two-stage calibration process described in Appendix H1., we implement several calibration alternatives, including (1) initial calibration, which uses the calibration results from the first stage without further fine-tuning; (2) human annotation, where a human operator utilizes a real-time GUI program to manually adjust the calibration parameters to align the rendered AirExo-2 contour with the visually observed contour from camera frames; and (3) two-stage calibration (mask only), which fine-tunes the calibration results by using only mask differences as supervision.
[…]
| Human Annotation | ||
| Two-Stage Calibration (mask only) | ||
| Two-Stage Calibration (mask + depth) | ||
[…]
Our two-stage calibration process achieves more accurate results compared to other alternatives. As reported in Tab. H1., our two-stage calibration process achieves the lowest error rates, with a 0.78% mask difference and 14.0 mm depth difference, yielding more precise calibration. It’s worth noticing that the depth difference here does not fully represent action accuracy because commercial depth sensors can produce noisy depth maps. Please refer to §LABEL:sec:accuracy-anaylsis for more details about action accuracy. Interestingly, human annotation performs even worse than initial calibration, mainly because annotators can only rely on 2D visual information to adjust calibration parameters. This limitation makes it difficult to accurately estimate depth information, leading to larger errors. Conversely, our two-stage calibration process explicitly models 3D information via differentiable rendering, offering a more reliable and precise solution for calibrating the AirExo-2 system.
[…]
Semi-Automatic SAM-2 Annotations. In the image adaptor, we initially annotate the hand mask (and, if visible, the head mask) manually using SAM-2 [sam2]. However, after a few annotations, we can fine-tune SAM-2 on the human-annotated samples, enabling automated labeling. This significantly reduces human effort and streamlines the demonstration adaptor process, making it nearly fully automated. ControlNet Training. We train a ControlNet [controlnet] based on the Stable Diffusion 1.5 [stable_diffusion] model to generate photo-realistic robot images from rendered robot images. To collect training samples, we use teleoperation to gather a small amount of play data, where the robot is teleoperated to move randomly within an empty workspace while recording RGB-D images and corresponding joint states. This ensures a diverse dataset of robot arm configurations, free from occlusions or distractions. Notably, these training samples are platform-specific but task-invariant, meaning they only need to be collected once per robot platform, and the trained ControlNet can be used across all tasks. This also opens up the possibility of directly transforming our in-the-wild demonstrations to other robotic arms without the need to design new exoskeletons that match specific robots. For training, we use a batch size of 88 and a learning rate of , while keeping other hyperparameters at their default settings. We use 50 DDPM sampling steps [ddpm] with a guidance scale of 9.0 [cfg]. The prompt for generating robot images for our robot platform is set to: robotic arms, dual arm, industrial robotic manipulator, metallic silver color, mechanical joints, precise mechanical details, gripper end effector, high-quality photo, photorealistic, clear and sharp details
[…]
[width=]images/dino_feat.pdf
Data Processing. The color image is resized to 448252 for DINOv2 backbone [dino] and 640360 for ResNet-18 backbone [resnet]. The depth image is resized to 640360 before creating the point cloud. The camera intrinsics are adjusted accordingly. Both the point clouds and actions are in the camera coordinate system. The point cloud is down-sampled with a voxel size of 5mm. For the data collected with teleoperation, we crop the point clouds using the range of [-0.7m, 0.7m], [-0.3m, 0.55m] and [0.9m, 1.55m]. For the data collected with AirExo-2, we crop the point clouds using the range of [-0.7m, 0.7m], [-0.3m, 0.45m] and [0.75m, 1.4m]. The robot trajectories are sampled using differences of translation, rotation and gripper width to remove redundant actions. For the action at two adjacent timesteps, if all the differences are less than the thresholds, only the first action is retained. The threshold for translation and gripper width is 5mm and the rotation threshold is . Network. The sparse encoder adopts a ResNet-like architecture built upon MinkowskiEngine [minkowski]. The dense encoder adopts DINOv2-base [dino] as the 2D backbone with the output channel of 128. In the spatial aligner, we use for feature alignment. The aligned features are fused by shared MLPs with the size of (256, 256, 256), and then fed into another sparse network. The two sparse networks are detailed in Tab. H1.. The transformer in action generator contains 4 blocks, in which we set 512 and 2048. The channel number of the readout token is 512. The diffusion head adopts a CNN implementation [dp] with 100 denoising iterations in training and 20 iterations in inference. The output action horizon used in experiments is 20. Training. RISE-2 is trained on 4 Nvidia A100 GPUs. The batch size is 240, the initial learning rate is 3e-4, and the warmup step is 2000. We employ a cosine scheduler to adjust the learning rate during training. 20% of the color images are augmented using a color jitter with (brightness, contrast, saturation, hue) parameters set to (0.4, 0.4, 0.2, 0.1).
[…]
| 2x mean pooling | ||
[…]
We conduct additional experiments on action representations for the Collect Toys task, comparing relative and absolute action representations. The results in Tab. H1. show that while relative action representation sometimes yields better results, absolute action representation provides more stable performance, particularly in terms of generalization. Therefore, we use absolute action representations throughout our experiments, except for the case study in §LABEL:sec:case-study.
[…]
| absolute | 40.0 | 12.5 | 5.0 | |
[…]
| absolute | 65.0 | 45.0 | 42.5 | |
[…]
To evaluate action accuracy, we designed a special evaluation board with three tracks, as illustrated in Fig. H1.. Each track has fixed holes spaced 2 cm apart. We created custom connectors for both the AirExo-2 and UMI [umi] that fit into these fixed holes, allowing us to collect position data. By sequentially placing the connectors into each fixed hole along the track, we can calculate the relative movement distance between two adjacent fixed holes and compare it with the true value (20 mm) to calculate the error.
[width=0.7]images/track.pdf
[…]
Fig. H1. visualizes the sparse semantic features obtained from the dense encoder by projecting the 2D feature map to 3D form using the reference coordinates. The sparse semantic features are aligned to the input point cloud using weighted spatial interpolation detailed in §LABEL:sec:rise2-arch. Although the 2D feature map output by the dense encoder is in low resolution (3218), we still observe clear and distinguishable continuous feature variations on the aligned features, where the targets at the current step can be easily identified from the entire scene. Such characteristic ensures precise feature fusion in the spatial domain. Additionally, we find that the features from DINOv2 change significantly as the task progresses, enabling the model to clearly understand the global state at the current time.